Becca is stringing a 28-foot floral garland along the 2 sides and top of a rectangular pergola to prepare for a wedding. The height of the pergola is 4 feet less than the width. Find the height and width of the pergola.

Answer by ikleyn(52788)

.
Height + Width + Height = 28


H      + W     + H      = 28


(W-4)  + W     + (W-4)  = 28


3W - 8 = 28


3W = 28 + 8 = 36


W = 12


ANSWER.  Width is 12 ft;  Height is  12-4 = 8 ft.

Solved.